# C-MELT: Contrastive Enhanced Masked Auto-Encoders for ECG-Language Pre-Training

### v2

@@ Abstract
Accurate interpretation of Electrocardiogram (ECG) signals is pivotal for diagnosing cardiovascular diseases. Integrating ECG signals with their accompanying textual reports holds immense potential to enhance clinical diagnostics through the combination of physiological data and qualitative insights. However, this integration faces significant challenges due to inherent modality disparities and the scarcity of labeled data for robust cross-modal learning. To address these obstacles, we propose C-MELT, a novel framework that pre-trains ECG and text data using a contrastive masked auto-encoder architecture. C-MELT uniquely combines the strengths of generative with enhanced discriminative capabilities to achieve robust cross-modal representations. This is accomplished through masked modality modeling, specialized loss functions, and an improved negative sampling strategy tailored for cross-modal alignment. Extensive experiments on five public datasets across diverse downstream tasks demonstrate that C-MELT significantly outperforms existing methods, achieving 15% and 2% increases in linear probing and zero-shot performance over state-of-the-art models, respectively. These results highlight the effectiveness of C-MELT, underscoring its potential to advance automated clinical diagnostics through multi-modal representations.
@@ 1 Introduction
Electrocardiograms (ECGs), obtained through non-invasive electrode placement, provide a critical window into the heart’s electrical activity by measuring voltage differences across specific anatomical regions. The standard 12-lead ECG, which captures unique electrical potential differences from each lead, plays a vital role in diagnosing a wide spectrum of cardiac conditions, like arrhythmias. In recent years, significant progress has been made in leveraging deep learning techniques for automated ECG interpretation (Yan et al., 2019; Ebrahimi et al., 2020; Siontis et al., 2021). However, these supervised deep learning approaches often necessitate large volumes of expertly annotated data, which are frequently scarce and expensive to acquire. Self-supervised learning (SSL) has emerged as a compelling alternative, offering the potential to learn robust representations from abundant unlabeled ECG data. These learned representations can be effectively utilized for zero-shot learning on novel tasks and adapted via fine-tuning to specific downstream applications, thereby mitigating the reliance on extensive labeled datasets.
@@
Numerous studies have explored the potential of SSL in the ECG domain, demonstrating its efficacy in learning representations from vast quantities of unlabeled data. These efforts generally fall into two main tracks: contrastive and generative approaches. Contrastive methods, exemplified by works such as (Chen et al., 2020; 2021; Chen & He, 2021; Grill et al., 2020; Kiyasseh et al., 2021; Oh et al., 2022; McKeen et al., 2024), aim to learn discriminative representations by maximizing the similarity between positive pairs (e.g., different augmentations of the same ECG signal) and minimizing the similarity between negative pairs (e.g., ECGs from different patients) within the embedding space. Conversely, generative approaches (Hu et al., 2023; Zhang et al., 2022a; 2023) focus on reconstructing the input data, typically by predicting masked or missing segments of the ECG signal, thereby learning to capture the underlying data distribution.
@@
Despite these advancements, existing ECG-based SSL approaches have largely overlooked the valuable information embedded within clinical text reports, which offer key insights into underlying cardiac conditions and have the potential to significantly enhance a model’s diagnostic accuracy (Zhang et al., 2022c; Chen et al., 2022). This oversight highlights a critical gap in the field: the lack of emphasis on jointly learning ECG-text cross-modal representations. While some recent efforts (Liu et al., 2024; Lalam et al., 2023; Li et al., 2024) have attempted to bridge this gap by integrating ECG signals and clinical reports through cross-modal contrastive learning, the potential of learning unified representations that capture the intricate interplay between ECG signals and their corresponding textual descriptions shown in generative approaches remains largely unexplored.
Moreover, the prevailing reliance on contrastive methods presents inherent limitations. These methods depend on the availability of negative samples and often struggle to capture cross-modal relationships effectively due to difficulties in defining appropriate negative pairings across different modalities (Liu et al., 2024). In contrast, generative approaches, such as (Na et al., 2024), hold significant promise for jointly learning robust ECG-text representations. By directly modeling the underlying data distribution across modalities, they can overcome some of the limitations inherent in contrastive methods (Kim et al., 2021). Therefore, integrating both contrastive and generative approaches within a unified framework could leverage their complementary strengths, leading to a more powerful method for learning robust cross-modal ECG-text representations.
@@
In this work, we depart from the reliance on either solely contrastive learning or stand-alone generative approaches for cross-modal representation learning. We introduce C-MELT, a novel hybrid framework that synergistically integrates both learning paradigms to effectively capture fine-grained input details and discriminative ECG-text features. Our approach employs tailored loss functions that promote the joint learning of robust cross-modal representations. Additionally, we introduce a nearest-neighbor negative sampling strategy, a crucial refinement often overlooked in previous methods, to ensure that negative samples are contextually relevant and challenging. Our novel sampling strategy demonstrably enhances the discriminative capacity of the learned representations, leading to improved performance across a variety of downstream tasks.
To rigorously evaluate the efficacy of C-MELT’s pre-trained encoders, we conduct extensive experiments on public ECG datasets in three evaluation settings: (1) Full fine-tuning: fine-tune the entire pre-trained ECG encoder with an added linear layer, adapting to specific downstream tasks while leveraging rich pre-learned representations. (2) Linear probing: freeze the pre-trained ECG encoder and train a linear classifier on top, directly assessing the quality and transferability of the learned representations with no adaptation. (3) Zero-shot learning: use the pre-trained ECG and text encoders to infer task-specific outputs without fine-tuning, relying solely on the learned representations to generalize to unseen tasks. Our experiments demonstrate that our method significantly outperforms recent state-of-the-art baselines across all evaluation settings and datasets.
@@ 2 Related Work
Self-supervised learning (SSL) has been shown to work effectively across various modalities, including vision (Li et al., 2022; Han et al., 2021), language (Devlin, 2018; He et al., 2020; Chung et al., 2024), and time-series data (Tonekaboni et al., 2021; Zhang et al., 2022b; Saeed et al., 2019). Particularly, recent advances in applying SSL to ECG signals have demonstrated that models can learn meaningful representations from large amounts of unlabeled data, which is crucial in medical domains where labeled datasets are often limited and expensive to acquire. Here, we mainly discuss two common SSL approaches: generative and contrastive, which have seen notable progress in ECG representation learning in recent years.
Early contrastive methods such as SimCLR (Chen et al., 2020), MoCo (Chen et al., 2021), SimSiam (Chen & He, 2021), and BYOL (Grill et al., 2020) introduced the concept of maximizing agreement between augmented views of the same data sample by employing augmentation strategies to create challenging positive and negative pairs. In the context of ECG signals, recent approaches like 3KG (Gopal et al., 2021) apply physiologically inspired spatial and temporal augmentations, using vectorcardiogram (VCG) transformations to capture the three-dimensional spatiotemporal characteristics of the heart’s electrical activity. Similarly, CLOCS (Kiyasseh et al., 2021) developed Contrastive Multi-Segment Coding (CMSC), which enhances the model’s ability to handle varying ECG signal characteristics across different axes—space, time, and patients. Building on this, (Oh et al., 2022) incorporates Wav2vec 2.0 (Baevski et al., 2020), CMSC, and random lead masking to simulate different global and local lead configurations during training, thereby improving model robustness and achieving impressive results on ECG downstream tasks.
@@
Multi-modal representation learning combines information from different data types, shown to effectively improve model performance (Lin et al., 2024; Du et al., 2023). Particularly, pioneering works like CLIP-based models (Radford et al., 2021; Rasheed et al., 2023; Zhai et al., 2023) have proven the power of contrastive learning in aligning visual and textual modalities, achieving strong generalization a broad range of tasks. Applying similar ideas to the ECG domain, MERL (Liu et al., 2024) leverages cross-modal and uni-modal alignment techniques to generalize ECG and text-based medical classification tasks. However, it overlooks the critical role of negative sample selection for contrastive learning and lacks exploring generative approaches for fine-grained multi-modal learning, limiting performance in end tasks.
@@ 3 Method
We propose C-MELT, a framework designed to learn generalizable cross-modal representations by aligning electrocardiogram (ECG) signals and corresponding medical text reports. C-MELT leverages masked language modeling (MLM) and masked ECG modeling (MEM) to reconstruct randomly masked segments within the input text and ECG signals, respectively. This encourages the model to learn fine-grained features within each modality. Furthermore, we introduce a contrastive objective based on the Siglep loss (Zhai et al., 2023) and a nearest-neighbor negative sampling strategy. Our objective directly promotes discriminative representation learning and enhances cross-modal alignment, besides the traditional ECG-text matching (ETM) learning task.
Figure 1 depicts the overall architecture of C-MELT, which comprises two main branches: an ECG encoder and a text encoder. The ECG encoder utilizes a transformer-based architecture (Vaswani et al., 2023) to process the input ECG signals and generate corresponding representations, denoted as $\mathbf{H}_{x}\in\mathbb{R}^{L_{x}\times d}$, where $L_{x}$ represents the sequence length of the ECG signal and $d$ represents the embedding dimension. The text encoder utilizes the recent pre-trained Flan-T5 model (Chung et al., 2024) which, to our knowledge, has not been previously applied to this task, to extract high-level semantic embeddings from the clinical text, denoted as $\mathbf{H}_{t}\in\mathbb{R}^{L_{t}\times d}$, where $L_{t}$ represents the sequence length of the text. These encoder outputs are then passed through a fusion module, which employs a cross-attention mechanism to integrate information from both modalities, generating fused representations denoted as $\mathbf{H}_{f}\in\mathbf{R}^{(L_{x}+L_{t})\times d}$. The model subsequently employs three distinct heads: two decoders, responsible for reconstructing the masked portions of the ECG signal ($\hat{X}$) and text ($T_{m}$), respectively, and a contrastive prediction head for ECG-text matching. Additionally, we introduce two projection heads, $g_{x}$ and $g_{t}$, following the ECG and text encoders, respectively. These projection heads, along with the Siglep loss, facilitate learning discriminative representation between these modalities. The model is trained by jointly optimizing four loss functions: masked language modeling loss ($\mathcal{L}_{MLM}$), masked ECG modeling loss ($\mathcal{L}_{MEM}$), ECG-text matching loss ($\mathcal{L}_{ETM}$), and the Siglep loss ($\mathcal{L}_{Siglep}$). The subsequent subsections provide a detailed description of each component within the C-MELT framework.
@@
We implement the ECG encoder (denoted as $\mathcal{F}_{x}$) based on a transformer architecture, which was originally developed for efficiently processing sequential data in parallel (Vaswani et al., 2023). We first follow (Oh et al., 2022) to apply a masking strategy to the ECG input $\mathbf{X}\in\mathbb{R}^{L\times C}$ to encourage robust feature learning, where $L$ is the length of the signal and $C$ is the number of channels. We then pass the masked input into a series of convolutional layers, each followed by GELU activation functions and group normalization. The extracted features are subsequently projected into a $768$-dimensional space. Next, we employ eight transformer encoder layers, each including a multi-head self-attention mechanism that allows the model to attend to different parts of the input sequence simultaneously. We conduct an experiment exploring the effects of different numbers of transformer layers in Section 4. Subsequently, we use a feed-forward network that projects the features into a $768$-dimensional space. Finally, we add a positional encoding layer to preserve the temporal order of the ECG sequence.
Language (Text) Encoder.
For our text encoder, we utilize the Flan-T5-base encoder (denoted as $\mathcal{F}_{t}$), which outputs 768-dimensional embeddings. The input to the encoder consists of token indices generated by the Flan-T5 tokenizer, represented as $\mathbf{T}\in\mathbb{Z}^{M}$, where $M$ is the maximum sequence length. Flan-T5 is an advanced version of the T5 model (Raffel et al., 2023), which has been pre-trained on a massive and diverse text dataset covering numerous tasks, such as summarization and question answering. We also conduct an ablation with various text encoders in Section 4.
@@
Decoders and Loss Functions.
After the fusion module, three distinct network heads are introduced, each associated with a specific loss function: masked language modeling (MLM), masked ECG modeling (MEM), and ECG-text matching (ETM). MLM and MEM are designed for reconstruction tasks, while ETM adopts a contrastive learning approach to align the different modalities. We detail each head and its corresponding loss function below:
Masked Language Modeling (MLM). The MLM head consists of a dense layer that outputs a probability distribution over the vocabulary. The MLM head focuses on predicting the masked tokens in the input text sequence, encouraging the model to learn contextualized word embeddings through a reconstruction task. We use the cross-entropy (CE) loss for MLM, as shown in Equation 1:
@@
Masked ECG Modeling (MEM). Similar to MLM, the MEM head aims to reconstruct the masked ECG inputs. It consists of a linear embedding layer that maps the input sequence to a lower-dimensional space (384), followed by learnable mask tokens that represent the missing portions of the sequence. We apply positional encodings to preserve the temporal structure of the ECG data. Subsequently, we use a multi-layer transformer decoder to model the dependencies within the sequence. Finally, a linear projection layer outputs the predicted ECG features. We train the MEM head using the mean squared error (MSE) loss between the predicted ECG features $\hat{\mathbf{x}}_{i}$ and the ground truth ECG features $\mathbf{x}_{i}$, as shown in Equation 2:
@@
ECG-Text Matching (ETM). Finally, we use ETM to promote alignment between ECG signals and their corresponding text reports. This is formulated as a binary classification task, where the ETM head consists of a single dense layer that outputs a scalar $\hat{z}_{\mathbf{x}_{k},\mathbf{t}_{k}}$ representing the predicted probability. The ETM loss is defined as the binary cross-entropy loss:
=$\mathcal{L}{{}_{\text{ETM}}}=-\frac{1}{\mathcal{B}}\sum_{k=1}^{\mathcal{B}}\left[y_{k}\log\sigma(\hat{z}_{\mathbf{x}_{k},\mathbf{t}_{k}})+(1-y_{k})\log(1-\sigma(\hat{z}_{\mathbf{x}_{k},\mathbf{t}_{k}}))\right],$ (3)
@@
3.2 Improving Contrastive Learning
Siglep Loss Function.
In multi-modal masked autoencoder architectures (Chen et al., 2022), contrastive learning’s effectiveness can be limited by the inherent tension between the reconstruction-focused generative tasks of autoencoders and the discriminative nature of contrastive learning. While generative tasks focus on reconstructing masked inputs, contrastive learning aims to distinguish between different data pairs. This can hinder the model’s capability to learn discriminative features useful for downstream tasks, such as zero-shot inference or linear probing. Furthermore, while the ETM loss in such architectures can serve as a contrastive loss, it may not be sufficient for building a robust ECG encoder. The ETM module is primarily designed for binary classification based on fused features rather than directly enhancing the discriminative power of individual encoders. This limitation can restrict the model’s ability to produce high-quality multimodal embeddings.
Therefore, we propose strengthening contrastive learning in multi-modal masked auto-encoder architectures using Siglep loss function. Specifically, we adapt the Siglip approach (Zhai et al., 2023), originally proposed for text-image pairs, to the text-ECG domain (Formula 4). This approach avoids the costly global normalization of softmax-based contrastive losses by operating independently on each ECG-text pair, improving memory efficiency and scalability. We introduce two additional network heads to the ECG and text encoders, respectively. Each head consists of a pooling layer, a Tanh activation function, and a dense layer, enabling them to output $768$-dimensional embeddings (denoted as $\mathbf{x^{\prime}}_{i}\in\mathbb{R}^{768}$ for the $i^{th}$ ECG sample and $\mathbf{t^{\prime}}_{j}\in\mathbb{R}^{768}$ for the $j^{th}$ text report).
=$\mathcal{L}{{}_{\text{Siglep}}}=-\frac{1}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\sum_{j=1}^{\mathcal{B}}\log\left(\frac{1}{1+e^{-y_{ij}\mathbf{x^{\prime}}_{i}^{\top}\mathbf{t^{\prime}}_{j}}}\right),$ (4)
@@ 3.2 Improving Contrastive Learning
In contrastive learning, the selection of negative samples significantly impacts the training process (Xu et al., 2022). Conventional methods often employ random sampling, where negative text reports are chosen randomly to replace positive texts. However, this approach may lead to false negative selection, especially in medical datasets, where randomly chosen reports might share substantial similarities with the positive reports, hindering effective contrastive learning. This is discussed more in the Appendix A.2.
Figure 2: Example of ECG-Text pair (left) and its corresponding negative text samples (right).
Therefore, we propose nearest-neighbor negative sampling (N3S), which selects negative samples based on their dissimilarity in the Flan-T5’s feature space, ensuring they are sufficiently distinct from the positive samples while remaining semantically related to the domain. Specifically, we first utilize pre-trained Flan-T5 (small) to generate vector representations, denoted as $\mathbf{v}_{t}\in\mathbb{R}^{512}$, for each text report $t$ in the training dataset $\mathcal{D}_{train}$. These embeddings capture the semantic meaning of the reports. During training, for a given ECG and its corresponding positive text report $(x_{k},t_{k}^{+})$ in a randomly selected half of the training batch $\mathcal{B}$, the negative report $t_{k}^{-}$ is selected as one of the top $64$ largest cosine distance reports from the positive report’s embedding $\mathbf{v}_{t_{k}^{+}}$. This makes the negative samples to be both challenging and distinct for effective contrastive learning.
To efficiently perform this process, we employ FAISS (Facebook AI Similarity Search) (Douze et al., 2024), a high-performance library designed for indexing and searching large collections of dense vectors. FAISS allows us to apply the N3S technique to large-scale datasets in a computationally tractable manner. Figure 2 shows one example of an ECG-text pair with its potential negative texts in the training dataset.

### v3

# Boosting Masked ECG-Text Auto-Encoders as Discriminative Learners

@@ Abstract
The accurate interpretation of Electrocardiogram (ECG) signals is pivotal for diagnosing cardiovascular diseases. Integrating ECG signals with accompanying textual reports further holds immense potential to enhance clinical diagnostics by combining physiological data and qualitative insights. However, this integration faces significant challenges due to inherent modality disparities and the scarcity of labeled data for robust cross-modal learning. To address these obstacles, we propose D-BETA , a novel framework that pre-trains ECG and Text data using a contrastive masked Auto-encoder architecture. D-BETA  uniquely combines the strengths of generative with Boosted Discriminative capabilities to achieve robust cross-modal representations. This is accomplished through masked modality modeling, specialized loss functions, and an improved negative sampling strategy tailored for cross-modal alignment. Extensive experiments on five public datasets across diverse downstream tasks demonstrate that D-BETA  significantly outperforms existing methods, achieving an average AUC improvement of 15% in linear probing with only one percent of training data and 2% in zero-shot performance without requiring training data over state-of-the-art models. These results highlight the effectiveness of D-BETA , underscoring its potential to advance automated clinical diagnostics through multi-modal representations. Our sample code and checkpoint are made available at https://github.com/manhph2211/D-BETA.
Cardiovascular Diagnostics, ECG-Text Multi-modal Representation Learning, Contrastive Masked Auto-Encoders
@@ 1 Introduction
Electrocardiograms (ECGs), obtained through non-invasive electrode placement, provide a critical window into the heart’s electrical activity by measuring voltage differences across specific anatomical regions. The standard 12-lead ECG, which captures unique electrical potential differences from each lead, plays a vital role in diagnosing a wide spectrum of cardiac conditions (e.g. arrhythmias). In recent years, significant progress has been made in leveraging deep learning techniques for automated ECG interpretation (Yan et al., 2019; Ebrahimi et al., 2020; Siontis et al., 2021). However, these supervised deep learning approaches often necessitate large volumes of expertly annotated data, which are frequently scarce and expensive to acquire. Self-supervised learning (SSL) has emerged as a compelling alternative, offering the potential to learn robust representations from abundant unlabeled ECG data. These learned representations can be effectively utilized for zero-shot learning on novel tasks and adapted via fine-tuning to specific downstream applications, thereby mitigating the reliance on extensive labeled datasets.
@@
Numerous studies have explored the potential of SSL in the ECG domain, demonstrating its efficacy in learning representations from vast quantities of unlabeled data. These efforts generally fall into two main tracks: contrastive and generative approaches. Contrastive methods, exemplified by works such as (Chen et al., 2020, 2021; Chen & He, 2021; Grill et al., 2020; Kiyasseh et al., 2021; Oh et al., 2022; McKeen et al., 2024), aim to learn discriminative representations by maximizing the similarity between positive pairs (e.g., different augmentations of the same ECG signal) and minimizing the similarity between negative pairs (e.g., ECGs from different patients) within the embedding space. Conversely, generative approaches (Hu et al., 2023; Zhang et al., 2022a, 2023) focus on reconstructing the input data, typically by predicting masked or missing segments of the ECG signal, thereby learning to capture the underlying data distribution. Therefore, integrating both contrastive and generative approaches within a unified framework could leverage their complementary strengths, leading to a more powerful method for learning robust ECG-text representations (Kim et al., 2021; Li et al., 2022b; Song et al., 2024).
@@
Despite advancements, existing ECG-based SSL approaches have largely overlooked the valuable information embedded within clinical text reports, which offer key insights into underlying cardiac conditions and have the potential to significantly enhance a model’s diagnostic accuracy (Zhang et al., 2022c; Chen et al., 2022). This oversight highlights a critical gap in the field: the lack of emphasis on jointly learning ECG-text cross-modal representations. Some recent efforts (Liu et al., 2024b; Lalam et al., 2023; Li et al., 2024; Liu et al., 2024a; Yu et al., 2024) have attempted to bridge this gap by integrating ECG signals and clinical reports through cross-modal contrastive learning, mainly employing relatively standard encoder models (e.g., ResNet (He et al., 2016), Bert (Devlin, 2018)). This makes the potential of learning robust representations that capture the intricate interplay between ECG signals and their corresponding textual descriptions, shown in generative approaches remain largely unexplored. Moreover, the prevailing reliance on these contrastive methods presents inherent limitations. They depend on the availability of negative samples and often struggle to capture cross-modal relationships effectively due to difficulties in defining appropriate negative pairings across different modalities.
@@
In this work, we depart from relying solely on contrastive learning or stand-alone generative approaches for cross-modal representation learning. We introduce D-BETA , a novel hybrid framework that synergistically integrates both learning paradigms to effectively capture fine-grained input details and discriminative ECG-text features. Our approach employs a transformer-based encoder specifically for ECG signals and a well-pre-trained language model for clinical text encoder in a masked auto-encoder architecture, together with tailored loss functions that promote the joint learning of robust cross-modal representations. Additionally, we introduce a nearest-neighbor negative sampling strategy, a crucial refinement often overlooked in previous methods, to ensure that negative samples are contextually selected and thereby, enhance the discriminative capability of the learned representations. To rigorously evaluate the efficacy of D-BETA , we conduct extensive experiments on various public ECG datasets and demonstrate that our method significantly outperforms recent state-of-the-art baselines across all evaluation settings.
@@ 2 Related Work
Self-supervised learning (SSL) has been shown to work effectively across various modalities, including vision (Li et al., 2022a; Han et al., 2021), language (Devlin, 2018; He et al., 2020; Chung et al., 2024), and time-series data (Tonekaboni et al., 2021; Zhang et al., 2022b; Saeed et al., 2019). Particularly, recent advances in applying SSL to ECG signals have demonstrated that models can learn meaningful representations from large amounts of unlabeled data, which is crucial in medical domains where labeled datasets are often limited and expensive to acquire. Here, we discuss two common SSL approaches: generative and contrastive, which have seen notable progress in ECG representation learning in recent years.
Early contrastive methods such as SimCLR (Chen et al., 2020), MoCo (Chen et al., 2021), SimSiam (Chen & He, 2021), and BYOL (Grill et al., 2020) introduced the concept of maximizing agreement between augmented views of the same data sample by employing augmentation strategies to create challenging positive and negative pairs. In the context of ECG signals, recent approaches like 3KG (Gopal et al., 2021) apply physiologically inspired spatial and temporal augmentations, using vectorcardiogram (VCG) transformations to capture the three-dimensional spatiotemporal characteristics of the heart’s electrical activity. Similarly, CLOCS (Kiyasseh et al., 2021) developed Contrastive Multi-Segment Coding (CMSC), which enhances the model’s ability to handle varying ECG signal characteristics across different axes-space, time, and patients. Building on this, (Oh et al., 2022) incorporates Wav2vec 2.0 (Baevski et al., 2020), CMSC, and random lead masking to simulate different global and local lead configurations during training, thereby improving model robustness and achieving impressive results on ECG downstream tasks.
@@
Multi-modal representation learning combines information from different data types, shown to effectively improve model performance (Lin et al., 2024; Du et al., 2023). Particularly, pioneering works like CLIP-based models (Radford et al., 2021; Rasheed et al., 2023; Zhai et al., 2023) have proven the power of contrastive learning in aligning visual and textual modalities, achieving strong generalizations across a broad range of tasks. Applying similar ideas to the ECG domain, recent efforts (Liu et al., 2024b; Yu et al., 2024; Liu et al., 2024a; Lalam et al., 2023) show promising progress in the field. Among them, MERL (Liu et al., 2024b) leverages cross-modal and uni-modal alignment techniques together with test-time clinical knowledge enhancement, which notably generalizes ECG and text-based medical zero-shot classification tasks. However, they often utilize genetic architectures (e.g., ResNet ECG encoder, Bert-based text encoder) and especially overlook the critical role of negative sample selection for contrastive learning and lacks exploring generative approaches for fine-grained multi-modal learning, limiting performance in end tasks.
@@ 3 Method
We propose D-BETA , a framework designed to learn generalizable cross-modal representations by aligning ECG signals and corresponding medical text reports. D-BETA  leverages masked language modeling (MLM) and masked ECG modeling (MEM) to reconstruct randomly masked segments within the input text and ECG signals, respectively. This encourages the model to learn fine-grained features within each modality. Furthermore, we introduce ETS  (ECG-Text Sigmoid) loss, as inspired by SigLIP (Zhai et al., 2023), and a nearest-neighbor negative sampling strategy. These directly promote discriminative representation learning and enhance cross-modal alignment, besides the ECG-text matching (ETM) learning task.
Figure 1 depicts the overall architecture of D-BETA , which comprises two main branches. The ECG encoder utilizes a transformer-based architecture (Vaswani et al., 2023) to process the input ECG signals and generate corresponding representations, denoted as $\mathbf{H}_{x}\in\mathbb{R}^{L_{x}\times d}$, where $L_{x}$ represents the sequence length of the ECG signal and $d$ represents the embedding dimension. The text encoder utilizes the recent pre-trained Flan-T5 model (Chung et al., 2024) which, to our knowledge, has not been previously applied to this ECG domain, to extract high-level semantic embeddings from the clinical text, denoted as $\mathbf{H}_{t}\in\mathbb{R}^{L_{t}\times d}$, where $L_{t}$ represents the sequence length of the text. These encoder outputs are then passed through a fusion module, which employs a cross-attention mechanism to integrate information from both modalities, generating fused representations denoted as $\mathbf{H}_{f}\in\mathbf{R}^{(L_{x}+L_{t})\times d}$. The model subsequently employs three distinct heads: two decoders, responsible for reconstructing the masked portions of the ECG signal ($\hat{X}$) and text ($T_{m}$), respectively, and a contrastive prediction head for ECG-text matching. Additionally, we introduce two projection heads, $g_{x}$ and $g_{t}$, following the ECG and text encoders, respectively. These projection heads, along with the ETS  loss, facilitate learning discriminative representation between these modalities. The model is trained by jointly optimizing four loss functions: masked language modeling loss ($\mathcal{L}_{MLM}$), masked ECG modeling loss ($\mathcal{L}_{MEM}$), ECG-text matching loss ($\mathcal{L}_{ETM}$), and the ETS  loss ($\mathcal{L}_{ETS~{}}$). The subsequent subsections provide a detailed description of each component within the D-BETA  framework.
@@
We implement the ECG encoder (denoted as $\mathcal{F}_{x}$) based on a transformer architecture, which was originally developed for efficiently processing sequential data in parallel (Vaswani et al., 2023). We first follow (Oh et al., 2022) to apply a masking strategy to the ECG input $\mathbf{X}\in\mathbb{R}^{L\times C}$ to encourage robust feature learning, where $L$ is the length of the signal and $C$ is the number of channels. Specifically, we leverage random lead masking as an on-the-fly augmentation where each lead randomly masked with a probability of $p=0.5$ during pre-training. Furthermore, we use a dropout layer on the input with $p=0.1$ to enable masking modeling. We then pass the masked input into a series of convolutional layers, each followed by GELU activation functions and group normalization. The extracted features are subsequently projected into a $768$-dimensional space. Following that, we add a convolutional positional encoding layer to preserve the temporal order of the ECG sequence. Next, we employ eight transformer encoder layers, each including a multi-head self-attention mechanism that allows the model to attend to different parts of the input sequence simultaneously. We conduct an experiment exploring the effects of different numbers of transformer layers in Section 4.3.
Text Encoder.
For our text encoder, we utilize the Flan-T5-base encoder (denoted as $\mathcal{F}_{t}$), which outputs 768-dimensional embeddings. The input to the encoder consists of token indices generated by the Flan-T5 tokenizer, represented as $\mathbf{T}\in\mathbb{Z}^{M}$, where $M$ is the maximum sequence length. Flan-T5 is an advanced version of the T5 model (Raffel et al., 2023), which has been pre-trained on a massive and diverse text dataset covering numerous tasks, such as summarization and question answering. Note that our text encoder is fine-tuned during the pre-training stage. In our attempt to demonstrate the effectiveness of recent Flan-T5 in the ECG-domain-based context, we also conduct an ablation with various text encoders in Section 4.3.
@@
Decoders and Self-Supervised Tasks.
After the fusion module, three network heads are introduced, each associated with a specific task or loss function: masked language modeling (MLM), masked ECG modeling (MEM), and ECG-text matching (ETM). MLM and MEM are designed for reconstruction tasks, while ETM adopts a contrastive learning approach to align the different modalities. We detail each task and its corresponding loss function below:
Masked Language Modeling (MLM). The MLM task consists of a dense layer that outputs a probability distribution over the vocabulary. The MLM head focuses on predicting the masked tokens in the input text sequence, encouraging the model to learn contextualized word embeddings through a reconstruction task. We use the cross-entropy (CE) loss for MLM, as shown in Equation 1:
@@
Masked ECG Modeling (MEM). Similar to MLM, the MEM task aims to reconstruct the masked ECG inputs. It consists of a linear embedding layer that maps the input sequence to a lower-dimensional space (384), followed by learnable mask tokens that represent the missing portions of the sequence. We apply positional encoding to preserve the temporal structure of the ECG data. Subsequently, we use a multi-layer transformer decoder to model the dependencies within the sequence. Finally, a linear projection layer produces the predicted ECG signal ($\hat{\mathbf{x}}_{i}$). We train the MEM head using the mean squared error (MSE) between this predicted signal and the ground truth signal ($\mathbf{x}_{i}$), as shown in Equation 2:
@@
ECG-Text Matching (ETM). Finally, we use ETM to promote alignment between ECG signals and their corresponding text reports, which further supports the fused feature space learning, together with generative aspects from MLM and MEM. This is formulated as a binary classification task, where the ETM task’s head consists of a single dense layer that outputs a scalar $\hat{z}_{\mathbf{x}_{k},\mathbf{t}_{k}}$ representing the predicted probability. The ETM loss is defined as the binary cross-entropy loss:
=$\mathcal{L}{{}_{\text{ETM}}}=-\frac{1}{\mathcal{B}}\sum_{k=1}^{\mathcal{B}}\Big{[}$ $y_{k}\log\sigma(\hat{z}_{\mathbf{x}_{k},\mathbf{t}_{k}})$
$+(1-y_{k})\log(1-\sigma(\hat{z}_{\mathbf{x}_{k},\mathbf{t}_{k}}))\Big{]},$ (3)
@@
3.2 ECG-Text Discriminative Learners.
ETS  Loss Function.
In multi-modal masked auto-encoder architectures such as (Chen et al., 2022), contrastive learning’s effectiveness can be limited by the inherent tension between the reconstruction-focused generative tasks of autoencoders and the discriminative nature of contrastive learning. They are more biased for learning to reconstruct masked inputs in generative manners. This can hinder the model’s capability to learn discriminative features useful for downstream tasks, such as zero-shot inference or linear probing. Furthermore, although the ETM loss in such architectures can serve as a form of contrastive loss, it may not be sufficient for building a robust ECG encoder. Specifically, the ETM module is primarily designed for binary classification based on fused features rather than directly enhancing the discriminative power of individual encoders. This limitation can restrict the model’s ability to produce high-quality multimodal embeddings.
Figure 2: Illustration of N3S in selecting negative samples within Flan-T5 space.
Therefore, we propose strengthening discriminative aspect in multi-modal masked auto-encoder architectures using ETS  loss, as shown in formula 4. This approach avoids the costly global normalization of softmax-based contrastive losses by operating independently on each ECG-text pair (sigmoid-based), improving memory efficiency and scalability. We introduce two additional network heads to the ECG and text encoders, respectively. Each head consists of a pooling layer, a Tanh activation function, and a dense layer, enabling them to output $768$-dimensional embeddings (denoted as $\mathbf{x^{\prime}}_{i}\in\mathbb{R}^{768}$ for the $i^{th}$ ECG sample and $\mathbf{t^{\prime}}_{j}\in\mathbb{R}^{768}$ for the $j^{th}$ text report).
=$\mathcal{L}{{}_{\text{ETS~{}}}}=-\frac{1}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\sum_{j=1}^{\mathcal{B}}\log\left(\frac{1}{1+e^{-y_{ij}\mathbf{x^{\prime}}_{i}^{\top}\mathbf{t^{\prime}}_{j}}}\right),$ (4)
@@
In contrastive learning, the selection of negative samples significantly impacts the training process (Xu et al., 2022). Conventional methods often employ random sampling, where negative text reports are chosen randomly to replace positive texts. However, this approach may lead to false negative selection, especially in medical datasets, where randomly chosen reports might share substantial similarities with the positive reports, hindering effective contrastive learning. This is discussed further in the Appendix A.2.
Therefore, we propose nearest-neighbor negative sampling (N3S), which selects negative samples based on their dissimilarity in the Flan-T5’s feature space (Figure 2), ensuring they are sufficiently distinct from the positive samples while remaining semantically related to the domain. Specifically, we first utilize pre-trained Flan-T5 (small) to generate vector representations, denoted as $\mathbf{v}_{t}\in\mathbb{R}^{512}$, for each text report $t$ in the training dataset $\mathcal{D}_{train}$. These embeddings capture the semantic meaning of the reports. During training, for a given ECG and its corresponding positive text report $(x_{k},t_{k}^{+})$ in half of the training batch $\mathcal{B}$, the negative report $t_{k}^{-}$ is selected as one of the top $64$ largest cosine distance reports from the positive report’s embedding $\mathbf{v}_{t_{k}^{+}}$. As the training progresses with batches being updated randomly, the negative samples continually change, introducing variability while maintaining domain relevance.
To perform this process, we employ FAISS (Facebook AI Similarity Search) (Douze et al., 2024), a high-performance library designed for indexing and searching large collections of dense vectors. FAISS allows us to apply N3S to large-scale datasets in a computationally tractable manner.